Bailey bought some of blue balloons for his birthday, the balloons inflated at the initial pressure of 1.70 atm, at that pressure the gas inside of it takes up 4.25L of volume, if Bailey increases the pressure to 2.40 atm, what will be the new volume?


1
Expert's answer
2022-04-05T11:26:05-0400

"P1V1= P2V2"

1.7× 4.25= 2.4× V

V2= 7.225/2.4

V2= 3.01L